OpenStudy (anonymous):

How do you solve the area of the rectangle, shown is 180 square centimeters (cm). Find its dimensions. Use the equation: L(L + 3) = 180. (You may want to distribute first.) |_| L+3 <------------- BOX L

OpenStudy (anonymous):

The area of the rectangle is obtained by multiplying the height and width: L(L+3)=180 L^2 +3L = 180 L^2 +3l -180 = 0 Solve using quadratic formula: L=12 and L=-15. We can't have negative length, so take L=12. So the dimensions of the rectangle are 12 and 3x12=36.